a vertical ring of radius r  and radius R falls vertically
 
it is in contact with 2 vertical rails which are joined at the top
 
the rails are without friction and resistance
 
there is a horizontal uniform magnetic field of magniitude B to the plane of th e rings and rail
 
when the speed of ring is v the current in section PQ is
 
 
Ans 8Brv/R
 
Please give detailed solutioon rates assured


Grade:10

1 Answers

ROSHAN MUJEEB
3 years ago
UpperpartoftherringhasresistanceR/2andlowerparthasresistanceR/2andtheyareconnectedinparallel,thus,effectiveresistance=R/4i=ResistanceEMF​=R/4Bv(2r)​=R8Bvr​